# Leveraging Latent Vector Prediction for Localized Control in Image Generation via Diffusion Models

### v1

@@ Abstract
Diffusion models emerged as a leading approach in text-to-image generation, producing high-quality images from textual descriptions. However, attempting to achieve detailed control to get a desired image solely through text remains a laborious trial-and-error endeavor. Recent methods have introduced image-level controls alongside with text prompts, using prior images to extract conditional information such as edges, segmentation and depth maps. While effective, these methods apply conditions uniformly across the entire image, limiting localized control. In this paper, we propose a novel methodology to enable precise local control over user-defined regions of an image, while leaving to the diffusion model the task of autonomously generating the remaining areas according to the original prompt. Our approach introduces a new training framework that incorporates masking features and an additional loss term, which leverages the prediction of the initial latent vector at any diffusion step to enhance the correspondence between the current step and the final sample in the latent space. Extensive experiments demonstrate that our method effectively synthesizes high-quality images with controlled local conditions.
@@
Local Control, Edge Control, Conditional Diffusion Models, Latent Space
@@ 1 Introduction
In recent years, large-scale Diffusion models [dalle, latentdiff, scalablediff, rectifiedflow] have become the new paradigm for text-to-image tasks. However, textual descriptions often prove to be insufficient in conveying the desired level of detail in the generated image. Following the saying “a picture is worth a thousand words”, recent methods, such as T2I-Adapter and ControlNet [t2iadapter, controlnet] have been proposed to incorporate image-level spatial conditions such as edges, segmentations, and depth maps into text-to-image generation. These methods are designed to apply conditions across the entire image, obeying the entirety of the prior condition map. Although these advances have gained significant attention for enabling more controllable image synthesis, they typically rely on conditioning inputs that span the entire image. As a result, fine-grained, localized control remains challenging, since the model requires prior information not only for the region of interest but also for the background.
@@
For example, T2I-Adapter [t2iadapter] interprets the absence of edges as smooth or contourless areas, focusing exclusively on the region of interest while neglecting the rest of the image. This produces backgrounds that lack significant detail, often leaving them almost empty. In contrast, ControlNet [controlnet] generates more complete images, but struggles to adhere to the specified conditions within the region of interest. Instead of maintaining the desired structure, ControlNet frequently introduces additional edges within this area. This deviation highlights their limitations in localizing the conditions precisely, despite its ability to enrich the overall image context. These behaviours are further analyzed through qualitative and quantitative comparisons in section 4.4 to support these observations.
@@ 2 Related Work
Another challenge in textual prompting arises when multiple subjects are included in the prompt. Often, one subject gains disproportionate attention, leading the model to neglect the other. To address this issue, methods like MultiConcept [multiconcept] have been using cross-attention layers, to ensure a balanced representation of all subjects in the generated image. Other approaches, [fastcomposer, collage, gligen] improve control over the composition by linking specific subject embeddings to image-level positions using masks or bounding boxes into the attention maps. Although greater control can be achieved with the latter models, they often lack the ability to define fine-grained details within those areas.
In order to achieve more fine-grained control over the text-to-image synthesis, methods like T2IAdapter and ControlNet [t2iadapter, controlnet] propose manipulating the latent feature space by introducing image-level features through the training of specialized encoders. This approach enables for precise control over the structure and composition of generated images but lack the ability to locally control them.
Figure 2: Overview of the T2I-Adapter integration. The input condition $c_{f}$ is processed by the adapter network $F_{\text{AD}}(c_{f})$ to generate condition features $F^{c}=\{F^{c}_{1},F^{c}_{2},F^{c}_{3},F^{c}_{4}\}$, which are designed to match the spatial dimensions of the corresponding encoder feature maps $F^{\text{enc}}=\{F^{\text{enc}}_{1},F^{\text{enc}}_{2},F^{\text{enc}}_{3},F^{\text{enc}}_{4}\}$ produced by the UNet. At each encoder layer, the adapter features are fused with the UNet features via element-wise addition, enabling the integration of image-level conditioning information into the diffusion process.
Some methods have extended T2I-Adapter [t2iadapter] and ControlNet [controlnet] to improve structural or appearance alignment, but without enabling explicit localization control. ControlNet++ [controlnet++] introduces a reward model that evaluates how well the generated image matches the input condition, encouraging global consistency but lacking the ability to focus control on a specific region. CTRL-X [ctrl-x] combines structure and appearance guidance by integrating findings from IP-Adapter [ipadapter], yet both the reference and structural images influence the entire output, limiting flexibility in background generation. SmartControl [smart_control] proposes modulating ControlNet’s skip connections to resolve conflicts between condition and prompt, learning a spatially-varying control strength. However, control still applies across the whole image and cannot be explicitly localized. Lastly, Text-Driven Image Editing [text_driven_regions] identifies salient objects to anchor control, but localization is determined automatically rather than by the user, and the method is specialized for object manipulation, making background freedom infeasible.
@@
Figure 3: Overview of our training methodology at each step $t$ of the denoising process. Given the textual prompt $p$, the control condition $c_{f}$ and region of interest defined by the mask $M$, an image I is passed through $\mathcal{E}$ to obtain its latent representation $z_{0}$. At each training step, we apply Local Control (see Section 3.3) to the T2I-Adapter features. A timestep $t$ is then selected, and diffusion is performed on $z_{0}$ to generate the noisy input $z_{t}$. Latent variable $z_{t}$ jointly with $y$ and the T2IAdapter features are passed into the denoising network $U$, to predict $\epsilon_{\theta}$. We subsequently Predict the initial latent vector $\widehat{z}_{0}$ (see Section 3.4) using $\epsilon_{\theta}$ and $z_{t}$ to Enforce Similiarity (see Section 3.5) by applying a Sobel filter $\mathcal{S}$ on the original and predicted latents $z_{0}$, $\widehat{z}_{0}$ within the region of interest.
@@ 3.1 Preliminaries
Figure 4: Comparison between direct masking of the edge map and our feature-level masking approach. When the mask is applied directly to the edge map, the model cannot distinguish between unstructured regions (intended smooth areas) and unconditioned regions (where no guidance should be applied), resulting in a loss of control outside the masked area. In contrast, our approach applies the mask to the T2I-Adapter features, preserving three distinct zones: structured regions with edge information, unstructured regions intended to be smooth, and unconditioned regions outside the mask. This separation enables precise, localized control while allowing the model to generate coherent and diverse content beyond the ROI.
@@ 3.3 Local Control through Masking
The integration of these masked features follows the same mechanism as in the original T2I-Adapter. Formally, the features from the adapter are first masked, using a downsample version of the binary mask $M$ to match the spatial resolution of each layer, and then injected into the UNet feature maps $F^{\text{enc}}$ at each layer $i$. The updated feature maps are computed as (6). This ensures that conditioning is only applied in the intended regions, while leaving the rest of the image unconstrained by structural guidance.
@@ 3.5 Enforcing Similarity
Building on the $\widehat{z}_{0}$ prediction derived in the previous section, we aim to enforce edge alignment within the ROI, thereby improving the T2IAdapter’s localized structural guidance. However, since $z_{0}$ is a latent representation of the generated image rather than its structure, it is necessary to extract edge features in the latent space before applying any penalty. To achieve this, we apply a high-pass filter to both the predicted $\widehat{z}_{0}$ and the original $z_{0}$ and compare them to apply the correspondent penalty.
@@ 4.1 Dataset
To train the T2IAdapter we used the 600k subset of the LAION-Aesthetics V2 [laion] public datasets, comprised of 600K pairs of text-image samples. For our purposes, we pre-processed the dataset to remove any pairs where the image resolution was lower than 256 pixels in any dimension. The remaining pairs were passed through canny edge algorithm to get the edge content. For validation and test purposes, we reserved 15% of the dataset pairs for validation and 5% for testing. We selected this dataset for its high variety and diversity of concepts, as well as its prior use in T2IAdapter training, making it particularly suitable for evaluating the controllability and fine-grained manipulation capabilities of our method.
@@ 4.2 Evaluation Metrics
Figure 6: Comparison between T2I-Adapter, ControlNet, and our proposed model for two different images with prompts a) “A tavern in the cold winter night” and b) ”A landscape with a river, stones and some trees in oil painting”. The figure shows the original images from which edges were extracted, as well as the mask used to determine the ROI (the right half of the edges), which is used as input for all models. The “Regions” column displays the regions of the original edge map, where the bottom-right portion is set to zero, leaving only the upper-right with the original edges. The subsequent columns present the generated images and their corresponding edge maps for each model. The values of DMSEin and DMSEout are also shown for these examples, where (green areas) indicate edges inside the ROI, and (red areas) indicate edges outside of it
@@ 4.3 Implementation Details
Table 1: Quantitative evaluation of T2I-Adapter, ControlNet, and our proposed method on 5,000 samples under the conditions of the Quadrant-Control Experiment. Reported metrics include Fréchet Inception Distance (FID), CLIPScore, and DMSEin. Lower FID and DMSEin values indicate better image quality and greater fidelity to the controlled area, respectively, while higher CLIPScore values reflect better image-text alignment.
@@ 4.5 Ablation Studies
Table 3: Quantitative evaluation of our method under different high pass filters. Specifically Sobel and Laplace on 5,000 samples under the Quadrant-Control Experiment. Reported metrics include Fréchet Inception Distance (FID), CLIPScore, and DMSEin.
@@
Table 4: Generative behaviour inside the controlled area. Results of varying ($\lambda$) and learning rate ($lr$) after 10k iterations, evaluated under the Quadrant-Control Experiment. Only the mean values are considered as it give us the orientation for which method produces higher fidelity on average.

### v2

# Localized Control in Diffusion Models via Latent Vector Prediction

@@ Abstract
Diffusion models emerged as a leading approach in text-to-image generation, producing high-quality images from textual descriptions. However, attempting to achieve detailed control to get a desired image solely through text remains a laborious trial-and-error endeavor. Recent methods have introduced image-level controls alongside with text prompts, using prior images to extract conditional information such as edges, segmentation and depth maps. While effective, these methods apply conditions uniformly across the entire image, limiting localized control. In this paper, we propose a novel methodology to enable precise local control over user-defined regions of an image, while leaving to the diffusion model the task of autonomously generating the remaining areas according to the original prompt. Our approach introduces a new training framework that incorporates feature-level masking and a penalization loss term, which leverages the prediction of the initial latent vector at any diffusion step to enhance the correspondence between the current step and the final sample in the latent space. Extensive experiments demonstrate that our method effectively synthesizes high-quality images with controlled local conditions. The code will be available once the paper is published.
@@
diffusion models; image generation; localized control; latent vector prediction; conditional image synthesis; masking features
@@ 1 Introduction
In recent years, large-scale Diffusion models [dalle, imagen, latentdiff, scalablediff, rectifiedflow] have become the new paradigm for text-to-image tasks. However, textual descriptions often prove to be insufficient in conveying the desired level of detail in the generated image. Following the saying “a picture is worth a thousand words”, recent methods, such as Text-to-Image Adapter (T2I-Adapter) and ControlNet [t2iadapter, controlnet] have been proposed to incorporate image-level spatial conditions such as edges, segmentations, and depth maps into text-to-image generation. These methods are designed to apply conditions across the entire image, obeying the entirety of the prior condition map. Although these advances have gained significant attention for enabling more controllable image synthesis, they typically rely on conditioning inputs that span the entire image. As a result, fine-grained, localized control remains challenging, since the model requires prior information not only for the region of interest but also for the background.
@@
For example, T2I-Adapter [t2iadapter] interprets the absence of edges as smooth or contourless areas, focusing exclusively on the region of interest while neglecting the rest of the image. This produces backgrounds that lack significant detail, often leaving them almost empty. In contrast, ControlNet [controlnet] generates more complete images, but struggles to adhere to the specified conditions within the region of interest. Instead of maintaining the desired structure, ControlNet frequently introduces additional edges within this area. This deviation highlights its limitations in localizing the conditions precisely, despite its ability to enrich the overall image context. These behaviours are further analyzed through qualitative and quantitative comparisons in section 4.4 to support these observations.
@@ 2 Related Work
Another challenge in textual prompting arises when multiple subjects are included in the prompt. Often, one subject gains disproportionate attention, leading the model to neglect the other. To address this issue, methods like Multiple Concept Control [multiconcept] have been using cross-attention layers, to ensure a balanced representation of all subjects in the generated image. Other approaches, [fastcomposer, collage, gligen] improve control over the composition by linking specific subject embeddings to image-level positions using masks or bounding boxes into the attention maps. Although greater control can be achieved with the latter models, they often lack the ability to define fine-grained details within those areas.
In order to achieve more fine-grained control over the text-to-image synthesis, methods like T2I-Adapter and ControlNet [t2iadapter, controlnet] propose manipulating the latent feature space by introducing image-level features through the training of specialized encoders. This approach enables for precise control over the structure and composition of generated images but lacks the ability to locally control them.
Figure 2: Overview of the T2I-Adapter integration. The input condition $c_{f}$ is processed by the adapter network $F_{\text{AD}}(c_{f})$ to generate condition features $F^{c}=\left\{F^{c}_{1},F^{c}_{2},F^{c}_{3},F^{c}_{4}\right\}$, which are designed to match the spatial dimensions of the corresponding encoder feature maps $F^{\text{enc}}=\left\{F^{\text{enc}}_{1},F^{\text{enc}}_{2},F^{\text{enc}}_{3},F^{\text{enc}}_{4}\right\}$ produced by the UNet. At each encoder layer, the adapter features are fused with the UNet features via element-wise addition, enabling the integration of image-level conditioning information into the diffusion process.
Some methods have extended T2I-Adapter [t2iadapter] and ControlNet [controlnet] to improve structural or appearance alignment, but without enabling explicit localization control. ControlNet++ [controlnet++] introduces a reward model that evaluates how well the generated image matches the input condition, encouraging global consistency but lacking the ability to focus control on a specific region. CTRL-X [ctrl-x] combines structure and appearance guidance by integrating findings from Image-Prompt Adapter (IP-Adapter) [ipadapter], yet both the reference and structural images influence the entire output, limiting flexibility in background generation. SmartControl [smart_control] proposes modulating ControlNet’s skip connections to resolve conflicts between condition and prompt, learning a spatially-varying control strength. However, control still applies across the whole image and cannot be explicitly localized. Lastly, Text-Driven Image Editing [text_driven_regions] identifies salient objects to anchor control, but localization is determined automatically rather than by the user, and the method is specialized for object manipulation, making background freedom infeasible.
@@
Figure 3: Overview of our training methodology at each step $t$ of the denoising process. Given the textual prompt $p$, the control condition $c_{f}$ and region of interest defined by the mask $M$, an image I is passed through $\mathcal{E}$ to obtain its latent representation $z_{0}$. At each training step, we apply Local Control (see Section 3.3) to the T2I-Adapter features. A timestep $t$ is then selected, and diffusion is performed on $z_{0}$ to generate the noisy input $z_{t}$. Latent variable $z_{t}$ jointly with $y$ and the T2I-Adapter features are passed into the denoising network $U$, to predict $\epsilon_{\theta}$. We subsequently Predict the initial latent vector $\widehat{z}_{0}$ (see Section 3.4) using $\epsilon_{\theta}$ and $z_{t}$ to Enforce Similiarity (see Section 3.5) by applying a Sobel filter $\mathcal{S}$ on the original and predicted latents $z_{0}$, $\widehat{z}_{0}$ within the region of interest.
@@ 3.1 Preliminaries
Figure 4: Comparison between direct masking of the edge map and our feature-level masking approach. When the mask is applied directly to the edge map, the model cannot distinguish between unstructured regions (intended smooth areas) and unconditioned regions (where no guidance should be applied), resulting in a loss of control outside the masked area. In contrast, our approach applies the mask to the T2I-Adapter features, preserving three distinct zones: structured regions with edge information, unstructured regions intended to be smooth, and unconditioned regions outside the mask. This separation enables precise, localized control while allowing the model to generate coherent and diverse content beyond the masked area.
@@ 3.3 Local Control through Masking
The integration of these masked features follows the same mechanism as in the original T2I-Adapter. Formally, the features from the adapter are first masked, using a downsampled version of the binary mask $M$ to match the spatial resolution of each layer, and then injected into the UNet feature maps $F^{\text{enc}}$ at each layer $i$. The updated feature maps are computed as (6). This ensures that conditioning is only applied in the intended regions, while leaving the rest of the image unconstrained by structural guidance.
@@ 3.5 Enforcing Similarity
Building on the $\widehat{z}_{0}$ prediction derived in the previous section, we aim to enforce edge alignment within the ROI, thereby improving the T2I-Adapter’s localized structural guidance. However, since $z_{0}$ is a latent representation of the generated image rather than its structure, it is necessary to extract edge features in the latent space before applying any penalty. To achieve this, we apply a high-pass filter to both the predicted $\widehat{z}_{0}$ and the original $z_{0}$ and compare them to apply the corresponding penalty.
@@ 4.1 Dataset
To train the T2I-Adapter we used the 600k subset of the LAION-Aesthetics V2 [laion] public datasets, comprised of 600K pairs of text-image samples. For our purposes, we pre-processed the dataset to remove any pairs where the image resolution was lower than 256 pixels in any dimension. The remaining pairs were passed through the canny edge algorithm to get the edge content. For validation and test purposes, we reserved 15% of the dataset pairs for validation and 5% for testing. We selected this dataset for its high variety and diversity of concepts, as well as its prior use in T2I-Adapter training, making it particularly suitable for evaluating the controllability and fine-grained manipulation capabilities of our method.
@@ 4.2 Evaluation Metrics
Figure 6: Comparison between T2I-Adapter, ControlNet, and our proposed model for two different images with prompts a) “A tavern in the cold winter night” and b) ”A landscape with a river, stones and some trees in oil painting”. The figure shows the original images from which edges were extracted, as well as the mask used to determine the ROI (the right half of the edges), which is used as input for all models. The “Regions” column displays the regions of the original edge map, where the bottom-right portion is set to zero, leaving only the upper-right with the original edges. The subsequent columns present the generated images and their corresponding edge maps for each model. The values of DMSEin and DMSEout are also shown for these examples, where (green areas) indicate detected edges inside the ROI, and (red areas) indicate detected edges outside of it
@@ 4.3 Implementation Details
Table 1: Quantitative evaluation of T2I-Adapter, ControlNet, and our proposed method on 5,000 samples under the conditions of the Quadrant-Control Experiment. Reported metrics include FID, CLIPScore, and DMSEin. Lower FID and DMSEin values indicate better image quality and greater fidelity to the controlled area, respectively, while higher CLIPScore values reflect better image-text alignment.
@@ 4.5 Ablation Studies
Table 3: Quantitative evaluation of our method under different high pass filters (Sobel and Laplace) applied to 5,000 samples from the Quadrant-Control Experiment. Reported metrics include FID, CLIPScore, and $\text{DMSE}_{\text{in}}$.
@@
Table 4: Generative behaviour inside the controlled area. Results of varying ($\lambda$) and learning rate ($lr$) after 10k iterations, evaluated under the Quadrant-Control Experiment. Only the mean values are considered as they provide orientation for which method produces higher fidelity on average.